OpenStudy (anonymous):

Simple....But I decided to ask it anyway. xD The sum of two numbers is 110. Their difference is 28. What is the smallest of the two numbers? A. 56 B. 41 C. 69 D. 55

Parth (parthkohli):

System of equations. Nuff said.

OpenStudy (apoorvk):

Okay then, then post what answer you got 'anyway'.

OpenStudy (anonymous):

41

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow x + y = 110 }\) \(\Large \color{MidnightBlue}{\Rightarrow x - y = 28 }\) Eliminate y by adding.

OpenStudy (anonymous):

how

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 2x = 138 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = 69 }\) As x is bigger, then x - 28 = y....69 - 28 = 41

Parth (parthkohli):

Get it?

OpenStudy (anonymous):

i got 69 also

Parth (parthkohli):

41 is the answer.

Parth (parthkohli):

it's asking for the smaller number, right?

OpenStudy (anonymous):

Yes

Parth (parthkohli):

So, it is 41

OpenStudy (merchandize):

The sum of two numbers is 110; x + y = 110 Their difference is 28; x - y = 28 x = 1/2(110 + 28) = 69 y = 41 The smaller of the two numbers is 41.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

how did u get 41

Parth (parthkohli):

x = 69{The larger number} 69 - y = 28 As shown -y = -41 y = 41

OpenStudy (anonymous):

k now i get it